OpenStudy (anonymous):

If a is increased by 25 percent and b is decreased by 25 percent, the resulting numbers will be equal. What is the ratio of a to b?

OpenStudy (anonymous):

a + a*.25 = b - b*.25

OpenStudy (phi):

why don't you try it?

OpenStudy (anonymous):

so

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

so

OpenStudy (phi):

you want to find a/b

OpenStudy (anonymous):

.75/1.25?

OpenStudy (anonymous):

3/5?

OpenStudy (anonymous):

is that really it?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Are the PSAT questions meant to look hard?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

But they just require thinking skills

OpenStudy (phi):

psat questions are a range of difficulty.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

I will do the next question and then report the results back to you guys

OpenStudy (anonymous):

s

OpenStudy (anonymous):

\[1.25a=0.75b\]\[\frac{1.25a}{b}=\frac{0.75b}{b}\to \frac{1.25a}{b}=0.75\]\[\frac{\frac{1.25a}{b}}{1.25}=\frac{0.75}{1.25}\to \frac{ a}{b}=\frac{0.75}{1.25}\to \frac{ a}{b}=\frac{3}{5}\]